OpenStudy (anonymous):

What is the greatest common divisor of 123 and 321?

OpenStudy (skullpatrol):

"Divisor" means the same as "factor."

OpenStudy (mathmate):

Factorize each and find the common factors. We see that 3 is a factor for each number because the sum of digits of each number is divisible by 3. So start divide each number by three to start factoring: 123=3*41 321=3*107 Since 41 and 107 are both primes, the factoring process is terminated. Can you see the common factor/divisor now?

OpenStudy (anonymous):

or you can use Euclid's algorithm which is incredible!

OpenStudy (mathmate):

Yep, I started thinking of using it for 3 digit numbers, unless of course there are obvious factors! Good thinking!